OpenStudy (anonymous):

Find the equation of a line which passes through point (1, 7) and is perpendicular to the line formed by the equation -2x + -10y = -9. Express it in slope-intercept form.

OpenStudy (anonymous):

Put (1,7) in\[y -y _{1}=m(x -x _{1})\]

OpenStudy (anonymous):

dasd

OpenStudy (anonymous):

dasd

OpenStudy (anonymous):

By the way, what is dasd?

OpenStudy (anonymous):

slope is y=1/5x+0.9 slope =1/5 1/5 * x =-1 x=-5 slope 2:=-5 -5(x,y)(1,7) y-7 =-5 ---- x-1 y-7=-5x+5 y=-5x+12

OpenStudy (anonymous):

Slope of given equation is -1/5

OpenStudy (anonymous):

perpendicular should multiply to -1 -1/5 * x=-1